compute the surface area generated when the first quadrant portion of the curve x^2-4y=8 from x1=0 to x2= 2 is revolved about the y-axis


1
Expert's answer
2022-06-03T13:00:00-0400
"y=\dfrac{x^2}{4}-2"

"y'=\dfrac{x}{2}"

"S=2\pi\displaystyle\int_{0}^{2}x\sqrt{(y')^2+1}dx"

"=2\pi\displaystyle\int_{0}^{2}x\sqrt{(\dfrac{x}{2})^2+1}dx"

"\int x\sqrt{\dfrac{x^2}{4}+1}dx"

"u=\dfrac{x^2}{4}+1, du=\dfrac{x}{2}dx"

"\int x\sqrt{\dfrac{x^2}{4}+1}dx=\int 2\sqrt{u}du"

"=2(\dfrac{2}{3})u^{3/2}+C=\dfrac{4}{3}(\dfrac{x^2}{4}+1)^{3/2}+C"


"S=2\pi\displaystyle\int_{0}^{2}x\sqrt{(y')^2+1}dx"

"=2\pi\displaystyle\int_{0}^{2}x\sqrt{(\dfrac{x}{2})^2+1}dx"

"=2\pi[\dfrac{4}{3}(\dfrac{x^2}{4}+1)^{3/2}]\begin{matrix}
 2\\
 0
\end{matrix}"

"=\dfrac{8\pi}{3}(2\sqrt{2}-1)({units}^2)"